### v1

# MiTTenS: A Dataset for Evaluating Misgendering in Translation

Abstract
Misgendering is the act of referring to someone in a way that does not reflect their gender identity. Translation systems, including foundation models capable of translation, can produce errors that result in misgendering harms. To measure the extent of such potential harms when translating into and out of English, we introduce a dataset, MiTTenS [^1] https://github.com/google-research-datasets/mittens, covering 26 languages from a variety of language families and scripts, including several traditionally underpresented in digital resources. The dataset is constructed with handcrafted passages that target known failure patterns, longer synthetically generated passages, and natural passages sourced from multiple domains. We demonstrate the usefulness of the dataset by evaluating both dedicated neural machine translation systems and foundation models, and show that all systems exhibit errors resulting in misgendering harms, even in high resource languages.
MiTTenS: A Dataset for Evaluating Misgendering in Translation
[…]
1 Introduction
It is well documented that dedicated machine translation systems show forms of gender bias (see Savoldi et al., 2021, for an overview). Prior work has highlighted bias when translating from source passages where the meaning is fundamentally ambiguous, in both academic and commercial systems Vanmassenhove et al. (2018); Johnson (2018, 2020). Forms of bias have been demonstrated with carefully constructed unambiguous English passages Stanovsky et al. (2019), and with linguistic constructions targeting specific language pairs (Cho et al., 2019; Bentivogli et al., 2020; Alhafni et al., 2022; Singh, 2023a, b; Stella, 2021, i.a.).
Recent advances have enabled general-purpose foundation models with powerful multilingual capabilities including translation Ouyang et al. (2022); OpenAI et al. (2023); Chung et al. (2022); Anil et al. (2023); Gemini Team Google (2023). These models can be used as building blocks in a wide range of products and applications, foregrounding other work on gender bias in natural language processing more broadly (Sun et al., 2019; Costa-jussà, 2019; Stanczak and Augenstein, 2021, i.a.).
Figure 1: Dataset examples targeting passages where translation errors may produce misgendering harms. Gender is encoded unambiguously in the source language (blue), and misgendering is highlighted in red.
Evaluating foundation models raises new challenges of measurement validity, given the wide range of use and potential harms (Weidinger et al., 2023; Shelby et al., 2023). Skew in training data and measures of bias in underlying models may not be reliable predictors or measurements of potential harm in downstream usage Goldfarb-Tarrant et al. (2021); Blodgett et al. (2020, 2021). There also remain challenges in empirically measuring performance as systems rapidly improve Jun (2023); Krawczyk (2023), ensuring high quality of service as multilingual capabilities expand Akter et al. (2023); Yong et al. (2023) and measuring unintentional harms in new system designs Renduchintala et al. (2021); Costa-jussà et al. (2023).
In this work, we focus on measuring misgendering harms of both dedicated translation systems and foundation models. Within translation, misgendering occurs when a translation refers to a person in a way that does not reflect their gender identity. We focus specifically on misgendering harms over other potential harms Costa-jussà et al. (2023), and expand coverage of language families and scripts at different levels of digital representation Stanovsky et al. (2019).
Adapting evaluation methods to measure misgendering for foundation models presents a few challenges. First, language models are often trained on public internet datasets Yang et al. (2023); Anil et al. (2023) which can cause contamination and render evaluation sets mined from public data sources ineffective Kiela et al. (2021). Second, gender is encoded in different ways across languages, making it challenging to scale automated evaluation methods. Automated methods enable faster modeling iteration, but methods commonly used in translation evaluations (eg, BLEU, BLEURT) may fail to capture specific dimensions of harm like misgendering. Finally, the evolving and contested nature of sociocultural norms related to gender make general purpose benchmark methods challenging to develop, particularly for expressions of non-binary gender across linguistic and cultural contexts globally Dev et al. (2021); Lauscher et al. (2023); Hossain et al. (2023); Cao and Daumé III (2020); Keyes (2018).
To address these challenges, we introduce Misgendered in Translation Test Set (MiTTenS); a new dataset with 13 evaluation sets, including 26 languages. We address challenges with contamination by creating targeted synthetic datasets, releasing provenance of mined datasets, and marking dataset files with canaries Srivastava et al. (2023). We address challenges with evaluation methods through evaluation sets that precisely target specific error patterns, many of which can be scored automatically with simple heuristics. We additionally release evaluation sets for translating out of English, for use with human evaluation protocols similar to Anil et al. (2023). To address varying sociocultural norms, we include multiple evaluation sets and first focus on errors where the potential for harm is unambiguous. Finally, we demonstrate the usefulness of the dataset across a range of dedicated translation systems like NLLB Team et al. (2022), and foundation models like GPT4 OpenAI et al. (2023).
High Mid Low Very low
Arabic Finnish Amharic Assamese
Chinese Indonesian Bengali Bhojpuri
French Polish Czech Lingala
German Telugu Farsi Luganda
Hindi Turkish Maithili
Italian Thai Oromo
Japanese
Portuguese
Russian
Spanish
# 2,252 488 784 108
Table 1: Languages included, grouped by level of digital resources, together with the number of examples in each group for translation into and out of English.
2 Dataset
In order to precisely target different constructions and languages, and to enable fine-grained disaggregated evaluation, MiTTenS contains multiple evaluation sets. Evaluation sets target potential harm when translating into English (‘‘2en’’), or when translating from English into another language (‘‘2xx’’). To enable automated evaluation, all 2en evaluation sets are constructed so that the source language input contains only a single gendered entity. This enables automated scoring of English translation by scanning for the expression of grammatical gender in personal pronouns. Each data point contains around 1-10 sentences per source passage, and additionally includes a reference translation. We add canaries to the CSV file to enable contamination checks, following Srivastava et al. (2023). We additionally release a data card Pushkarna et al. (2022) on GitHub.
Eval set Subset #
2xx: Translating out of English
Gender Sets coref:coreference 592
Gender Sets coref:synthetic S 224
Gender Sets gender_agreement:contextual S 496
Gender Sets gender_agreement:news 192
Gender Sets gender_agreement:wiki 256
Gender Sets gender_specific S 128
2en: Translating into English
Gender Sets coref:coreference 180
Gender Sets coref:synthetic S 210
Gender Sets gender_agreement:contextual S 120
Gender Sets gender_specific S 120
Late binding late_binding 252
Enc in nouns nouns_then_pronouns 222
SynthBio synthbio S 640
Table 2: Datasets for measuring misgendering harms. S marks synthetic data, # marks number of examples.
2.1 Gender Sets
The Gender Sets evaluation set was built from error analysis in publicly available translation systems. The linguistic phenomena targeted include co-reference (Polish “ Mój przyjaciel jest piosenkarzem, ale kompletnie bez talentu” to English “My friend is a singer but he is not talented at all”), gender agreement (Spanish “Mario trabaja como empleado doméstico. Casi no pasa tiempo en su casa…” to English “Mario works as a housekeeper. He rarely spends time at home.”), and gender-specific words (English “I went to my mother’s house yesterday. She is British.” to French “Je suis allé chez ma mère hier. Elle est britannique.”).
Examples targeting co-reference were created using a mix of handwritten and synthetic methods. Examples targeting gender agreement were created from three sources: adapted from Translated Wikipedia Biographies Stella (2021), sourced from public news websites, or created synthetically. Examples targeting gender-specific words were created synthetically. Professional translators were used in creating reference translations. In total, this consists of 1,888 2xx data points. To enable automated evaluation for all 2en evaluation sets, we additionally filter those examples down to 630 2en data points. Filtering removes source passages with more than one English gender pronoun, and languages like Bengali that do not encode gender information in pronouns.
2.2 SynthBio
The SynthBio evaluation set is mined from a subset of Yuan et al. (2022), which includes synthetically generated English biography passages with multiple sentences. Using synthetic data avoids potential data contamination from sources like Translated Wikipedia Biographies Stella (2021), which language models may have seen during pre-training. We filter SynthBio to only include passages encoding a single gendered entity with binary pronouns, then take a stratified sample based on English gender pronouns, and finally create pairs for a subset of languages using machine translation. This consists of 640 examples targeting translation into English. These passages often require gender information to be translated correctly across multiple sentences, and are longer passages. An example Thai to English reference translation is:
Suzanne Abamu was a Congolese feminist theologian, professor, and activist. Abamu was born on April 12, 1933 in Dékolé, Republic of the Congo. She attended the University of Sorbonne Paris. She died on February 22, 2012 in Paris due to renal failure. She is buried in Cimetiere du Montparnasse in Paris. She is the daughter of Maria Abamu and Augustin Abamu. Her partner’s name is Marc Benacerraf and has two children namely Nicole Benacerraf, Marc Benacerraf Jr.
2.3 Late binding
The Late binding evaluation set was created from error analysis on translation errors in Gender Sets and targets passages in Spanish with where the gender information is only encoded later in the passage, but where an English translation would require expression of gender early in the translation. For example in Spanish “Vino de inmediato cuando se enteró porque es una buena bibliotecaria” does not encode gender information until the end of the sentence, but in an English translation gender information would come early in “She came right away when she found out because she is a good librarian.” This evaluation set uses a mix of nouns for family names as well as a subset of nouns from Winogender Rudinger et al. (2018), and consists of 252 examples targeting translation into English.
Figure 2: Evaluation results using automated evaluation when translating into English. Gemini and PaLM 2 systems perform best when considering worst-case performance, and GPT4 is within 5 percentage points.
2.4 Encoded in nouns
The Encoded in nouns evaluation set targets languages like Finnish that don’t encode gender information in personal pronouns but do encode gender information lexically through the choice of noun word (e.g., isä or äiti). This consists of 222 examples targeting translation into English. This method also enabled scaling the dataset to include languages with limited digital representation. An example from the evaluation set in Oromo is “Saaraan akkoo kooti. Qoosaa ishee baay’een jaalladha.” with a reference translation of “Sarah is my aunt. I really like her jokes.”
Overall Weakest Weakest Worst-case
Family Model accuracy language evaluation set performance
[…]
GPT 4 gpt-4-1106-preview 99.1% Lingala Enc in nouns 66.7%
GPT 3.5 gpt-3.5-turbo-1106 95.9% Amharic Late binding 42.9%
Gemini gemini-pro 97.8% Spanish Late binding 71.4%
PaLM 2 text-bison-001 99.0% Indonesian Late binding 71.4%
PaLM 2 text-bison-32k 98.4% Hindi Late binding 71.4%
Mistral Mistral-7B-Instruct-v0.1 92.7% Lingala Late binding 14.3%
[…]
3 Evaluation
MiTTenS can be used in evaluation for external audits of a deployed system, hill-climbing during model development, or monitoring during training. Here, we demonstrate using the dataset for automated evaluation of 2en translation with a range of systems. For an example 2xx human evaluation protocol see Anil et al. (2023). We leave demonstration of LLM-based evaluation Zheng et al. (2023); OpenAI et al. (2023) for future work.
We demonstrate evaluation using multiple evaluation sets, and report overall accuracy but also report performance when disaggregated by language, by evaluation set, and by worst-case performance across the combination of these dimensions. This approach scales as additional evaluation sets are added, and more precise evaluation data enables targeted interventions or improvements.
GPT systems were queried with the OpenAI Python client, and PaLM 2 and Gemini systems with the Cloud Vertex Python SDK. Mistral was evaluated through a HuggingFace Endpoint. NLLB was run in local inference. Foundation models were prompted with an instruction “Translate the following text from Turkish to English” with greedy sampling (top-k=1 or temperature=0).
Evaluation results are shown in Figure 2 and Table 3. Even though systems show relatively high overall accuracy, all systems perform worse on passages that require translation to “she” as compared to “he”, which may be related to patterns of representation in training datasets Chowdhery et al. (2022). Performance is often worst on Encoded in nouns or Late binding evaluation sets. Surprisingly, we see areas of weakness even in high resource languages, and different areas of weakness in the same model families. There is no clear pattern to which languages are most challenging across systems, demonstrating the importance of empirical evaluations, and that MiTTenS can be used to pinpoint areas for targeted improvement.
4 Conclusion
We release MiTTenS, a dataset for measuring misgendering harms with 13 evaluation sets that covers 26 languages. This dataset makes progress towards more precisely measuring potential harms and scaling evaluation to more languages. We address challenges with contamination and scoring methods amidst evolving sociocultural norms.
Future research should investigate how increasingly capable foundation models might enable products with multiple potential translations. More work is required to work towards language technologies that produce accurate and faithful representations of all people.
Limitations
For gender-related errors in translation systems, evaluations do not consider differential harms to people related to expressing non-binary gender identities (Keyes, 2018; Dev et al., 2021; Lauscher et al., 2023), or consider contested perspectives on pronouns across languages and cultures (Lee, 2019). Moreover, while gender agreement into English is amenable to automatic evaluation, evaluation of gender agreement out of English remains challenging and time-intensive. This dataset does not include examples for direct translation between languages beyond English. Finally, we note that our evaluations focus on only a subset of potential risks (Weidinger et al., 2021), and that our evaluations focus on model outputs without considering the wider sociotechnical context in which instruction-finetuned language models exist (Weidinger et al., 2023; Shelby et al., 2023).
Ethical Considerations
This work aims to contribute to society and to human well-being by creating new dataset and demonstrating how they can be used to avoid potential harms in translation systems. Improving the quality of measurement and evaluation is a critical aspect of building fair and inclusive translation systems. However, we also acknowledge that not all possible gender related harms and errors may have been covered in this work, and thus, it should not be used as a singularly used dataset to certify any translation system free of potential harm.
Acknowledgements
We thank Marie Pellat, Orhan Firat, Kellie Webster, Kathy Meier-Hellstern, Erin van Liemt, Mark Díaz, and Amber Ebinama for their input, feedback, and advice.